# Quasi-convex Splittings of Acylindrical Graphs of Locally finite-height Groups

### v1

@@ 3.3. Application to Hyperbolic One-Relator Groups
We now apply Theorem 3.3 to Question 1.3 via one-relator groups.
@@
Let $G$ be a one-relator group, and $\langle X\mid w\rangle$ a one-relator presentation for $G$. We say that a subgroup $H\leq G$ is a Magnus subgroup with respect to this presentation if there exists some subset $Y\subset X$ such that $w\notin\langle Y\rangle<F(X)$, and $Y$ generates $H$ in $G$. We say that a subgroup is a Magnus subgroup if it is a Magnus subgroup relative to some one-relator presentation of $G$.
@@
    Each $G_{i}$ is a one-relator group, and $G_{0}$ is isomorphic to a finite free product of cyclic groups, and;
@@
In the language of Definition 2.3, a Magnus hierarchy for a one-relator group $G$ is a hierarchy decomposition where $J_{i}=\{1\}$ for all $i$, each group $G_{i,1}$ is one-relator and is an HNN-extension of $G_{i-1,1}$ given by identifying Magnus subgroups, with the addition of one more level if necessary to break $G_{0}$ into its cyclic free factors.
@@
([Masters06], see also [MagnusSplitting, Moldavanskii]) Let $G$ be a one-relator group. Then $G$ admits a Magnus hierarchy.
Returning to acylindricity, we have the following powerful result due to Linton, where the definition of $\mathbb{Z}$-stability can be found in [Linton22, Section 1.3].
@@
([Linton22, Theorem 7.1]) Let $G$ be a one-relator group and $G_{0}\hookrightarrow G_{1}\hookrightarrow\cdots\hookrightarrow G_{n}\cong G$ a Magnus hierarchy for $G$. Then the following are equivalent.
@@
It may be shown by direct calculation that this presentation is $C^{\prime}(1/6)$, and thus $G$ is hyperbolic. Using the homomorphism $G\rightarrow\mathbb{Z}$ that sends $a$ to $1$ and $b$ to 00 we may express $G$ as the HNN-extension of the one-relator group
$G_{1}=\langle a,b,c,d\mid dc^{2}ad^{-2}ca^{-2}ba^{-1}b^{-1}\rangle$
along the subgroup isomorphism $\phi_{1}:\langle a,b,c\rangle\mapsto\langle b,c,d\rangle$, and in turn we may use the homomorphism $G_{1}\rightarrow\mathbb{Z}$ that sends $b$ to $1$ and all other generators to 00 to express $G_{1}$ as the HNN-extension of the free group
$G_{2}=\langle a,c,d,a_{+},c_{+},d_{+}\mid dc^{2}ad^{-2}ca^{-2}a_{+}\rangle\cong F(a,c,d,c_{+},d_{+})$
along the subgroup isomorphism $\phi_{2}:\langle a,c,d\rangle\mapsto\langle a_{+},c_{+},d_{+}\rangle$. This is therefore a Magnus hierarchy for $G$, which we claim is acylindrical. Indeed, one may verify by hand or using the pseudo-algorithm of Linton [LintonThesis, Proposition 6.1.7] that both extensions are $\mathbb{Z}$-stable, and the fact that $G$ is hyperbolic implies that it is Baumslag-Solitar free. Thus, by Theorem 3.8 this hierarchy is acylindrical, and is in turn a quasi-convex hierarchy for $G$.
Let $A$ be some edge group of the first splitting in this hierarchy, so by construction the normal closure of $A$ is the kernel of the homomorphism $G\rightarrow\mathbb{Z}$ that sends $a$ to $1$ and $b$ to 00. One may verify that this is not finitely generated using Brown’s algorithm [BrownBNS, Theorem 4.3]. However, the group $G$ does fibre — by the same algorithm the homomorphism $\chi:G\rightarrow\mathbb{Z}$ that sends $a$ to $1$ and $b$ to $-1$ has finitely generated kernel, which we will call $F$.
Consider now the intersection of $F$ with any $G$-conjugate $A^{g}$ of $A$. We must have that $A^{g}\not\subset F$ because the normal closure of $A$ is the infinitely generated kernel of a map to $\mathbb{Z}$, and so the quotient of $A^{g}$ by $F\cap A^{g}$ must be isomorphic to $\mathbb{Z}$. It follows that, since $A^{g}$ is free of rank $3$, $F\cap A^{g}$ cannot be finitely generated. The splitting of $F$ induced by the above hierarchy therefore does not have finitely generated edge groups, but will be acylindrical (and indeed finite-height by Theorem 3.8).
While the group $F$ in Example 3.9 is a non-elementary free group, so naturally admits a free splitting, the above argument illustrates that acylindrical graphs (or hierarchies) of free groups may not induce splittings with finitely generated edge groups on their subgroups. However, we may still deal with such subgroups using Theorem 3.3. We have the following corollary of Theorem 3.3, a reformulation of Corollary 1.6, which shows that those subgroups whose actions induced by a hierarchy do not have finitely generated edge stabilisers will nevertheless admit finite-height hierarchies with finitely generated edge stabilisers. Assuming Conjecture 1.2 this therefore answers Question 1.3 in the affirmative for all finitely generated subgroups of one-relator groups with acylindrical hierarchies.
@@
Let $H$ be a finitely generated subgroup of a one-relator group $G$ such that $G$ admits an acylindrical Magnus hierarchy $G_{0}\hookrightarrow G_{1}\hookrightarrow\cdots\hookrightarrow G_{n}\cong G$. Then $H$ admits a non-trivial hierarchy of non-trivial finite-height splittings with finitely generated edge stabilisers. In particular the group $H$ is acylindrically arboreal if and only if it admits a non-trivial finite-height splitting with finitely generated edge stabilisers.
@@
First, we observe that by [Coherence1r, Theorem 1.1] $H$ will be finitely presented. We may assume without loss of generality that $H$ is not a subgroup of $G_{n-1}$, and proceed by induction on $n$. First assume $n=0$. Then any finitely generated subgroup of $G$ is a free product of cyclic groups, and so naturally admits a finite-height splitting whose vertex groups are cyclic as required.
@@
Now assume that the above result holds for all subgroups of one-relator groups with a Magnus hierarchy of length at most $k-1$, and assume that $n=k$. By Theorem 3.8, $G$ is hyperbolic and the Magnus Hierarchy for $G$ is quasi-convex, so $G$ splits as an HNN-extension $G_{n-1}*_{F}$, where $F$ is a free subgroup of $G$ that is quasi-convex in $G$. This splitting is therefore finite-height by Theorem 2.19, and so acylindrical by Lemma 2.11 and the fact that $G$ is hyperbolic and thus has bounded finite subgroups. The group $H$ acts non-trivially on the Bass–Serre tree $T$ associated to this decomposition by the assumption that $H\not\leq G_{n-1}$, and every edge of the minimal subtree $T_{H}$ of $T$ associated to the $H$ and guaranteed by Lemma 2.1 is in the $G$-orbit of the edge of $T$ corresponding to $F$ which is locally finite-height and has finite-height in $G$ by construction. We may therefore invoke Corollary 3.4(b) to see that $H$ admits a finite-height splitting with finitely generated edge stabilisers, whose associated Bass–Serre tree we will call $T^{\prime}$, and which dominates the action of $H$ on $T_{H}$. The vertex stabilisers of the action of $H$ on $T^{\prime}$ will be subgroups of the vertex stabilisers of the action of $H$ on $T_{H}$, which in turn are conjugate into of $G_{n-1}$ in $G$ as they arise as stabilisers of the action of $H$ on $T$ induced by the action of $G$ on $T$. They are therefore all isomorphic to subgroups of $G_{n-1}$, a one-relator group with an acylindrical hierarchy of length $n-1=k-1$, and so admit hierarchies of non-trivial finite-height splittings with finitely generated edge stabilisers by the inductive hypothesis. It follows that $H$ must also admit such a hierarchy.
@@
Now assume that $H$ is acylindrically arboreal. It is therefore not virtually cyclic by Theorem 2.6, and so the hierarchy constructed in the first part of this proof will not be trivial unless $H\leq G_{0}$, and is a free product of cyclic groups. Taking the first stage of this hierarchy therefore gives the required non-trivial finite-height splitting with finitely generated edge stabilisers.
@@ 3.4. Relaxing The Conditions on Stabilisers
The group $G_{1}$ is recursively presented, so there exists a finitely presented group $G_{2}$ such that $G_{1}\leq G_{2}$ is a malnormal subgroup by a result of Wagner [Wagner24, Theorem 1.1]. Define $G$ to be the trivial HNN-extension of $G_{2}$ along $\langle t_{0}\rangle$, so $G=\langle G_{2},t\mid[t,t_{0}]\rangle$.
@@
$g\cdot\phi(\tilde{v})=\phi(g\cdot\tilde{v})=\phi(\tilde{v}),$
so $(H_{1}*_{\langle t_{0}\rangle}\mathbb{Z}^{2})\leq\operatorname{PStab}_{G}(\phi(\tilde{v}))$. By definition of $T$ this will hold only when $\phi(\tilde{v})=v$. It follows that $(H_{1}*_{\langle t_{0}\rangle}\mathbb{Z}^{2})\leq\operatorname{PStab}_{G}(\tilde{v})\leq(H_{1}*_{\langle t_{0}\rangle}\mathbb{Z}^{2})$, so $\operatorname{PStab}_{G}(\tilde{v})=(H_{1}*_{\langle t_{0}\rangle}\mathbb{Z}^{2})$ as required, and that at least one vertex stabiliser of the action of $G$ on $T^{\prime}$ is infinitely generated. The result then follows by Lemma 2.2. ∎
@@
The Dunwoody construction applied to the acylindrical splitting of the group $G$ above will not preserve acylindricity.
@@
which is acylindrical due to the malnormality of $G_{1}$ in $G_{2}$. We do however believe that there will be an example of a finitely presented group that acts acylindrically on a tree but not with finitely generated edge stabilisers, and finish this paper by restating Question 1.9.
Question 3.13.

### v2

@@ 3.3. Application to Hyperbolic One-Relator Groups
We now apply the methods used in the proof of Theorem 3.3 to Question 1.3 via one-relator groups.
@@
Let $G$ be a one-relator group, and $\langle X\mid w\rangle$ a one-relator presentation for $G$ with $w$ cyclically reduced as a word over $X$. We say that a subgroup $H\leq G$ is a Magnus subgroup with respect to this presentation if there exists some subset $Y\subset X$ such that $w\notin\langle Y\rangle<F(X)$, and $Y$ generates $H$ in $G$. We say that a subgroup is a Magnus subgroup if it is a Magnus subgroup relative to some one-relator presentation of $G$.
@@
    Each $G_{i}$ is a one-relator group, and $G_{0}$ is isomorphic to the free product of finitely many cyclic groups, and;
@@
In the language of Definition 2.3, a Magnus hierarchy for a one-relator group $G$ can be identified with a hierarchy decomposition of length $n+1$:
  • For all $i\neq 0$, set $J_{i}=\{1\}$ and let $G_{i,1}\cong G_{i-1}$. Then for $i\geq 2$, $T_{i,1}$ may be chosen to be the $HNN$-extension decomposition of $G_{i-1}$ with edge stabilisers isomorphic to $G_{i-1}$ coming from the Magnus hierarchy.
  • The group $G_{1,1}\cong G_{0}$ is the free product of finitely many cyclic groups, so acts on a tree with finite vertex stabilisers by [undefs]. We may therefore choose $T_{1,1}$ to be this tree to satisfy the condition in Definition 2.3 that all groups $G_{0,j}$ for $j\in J_{0}$ must be finite.
@@
([undefx], see also [undefw, undefac]) Let $G$ be a one-relator group. Then $G$ admits a Magnus hierarchy.
Returning to acylindricity, we have the following powerful result due to Linton, where the definition of $\mathbb{Z}$-stability can be found in [undefu, Section 1.3].
@@
([undefu, Theorem 7.1]) Let $G$ be a one-relator group and $G_{0}\hookrightarrow G_{1}\hookrightarrow\cdots\hookrightarrow G_{n}\cong G$ a Magnus hierarchy for $G$. Then the following are equivalent.
@@
It may be shown by direct calculation that this presentation is $C^{\prime}(1/6)$, and thus $G$ is hyperbolic. Using the homomorphism $G\rightarrow\mathbb{Z}$ that sends $a$ to $1$ and $b$ to 0 we may express $G$ as the HNN-extension of the one-relator group
$G_{1}=\langle c,d,e,f\mid fe^{2}cf^{-2}ec^{-2}dc^{-1}d^{-1}\rangle$
along the subgroup isomorphism $\phi_{1}:\langle c,d,e\rangle\mapsto\langle d,e,f\rangle$, and in turn we may use the homomorphism $G_{1}\rightarrow\mathbb{Z}$ that sends $b$ to $1$ and all other generators to 0 to express $G_{1}$ as the HNN-extension of the free group
$G_{2}=\langle g,h,i,k,l,m\mid ih^{2}gi^{-2}hg^{-2}k\rangle\cong F(g,h,i,l,m)$
along the subgroup isomorphism $\phi_{2}:\langle g,h,i\rangle\mapsto\langle k,l,m\rangle$. This is therefore a Magnus hierarchy for $G$, which we claim is acylindrical. Indeed, one may verify by hand or using the pseudo-algorithm of Linton [undeft, Proposition 6.1.7] that both extensions are $\mathbb{Z}$-stable, and the fact that $G$ is hyperbolic implies that it is Baumslag-Solitar free. Thus, by Theorem 3.8 this hierarchy is acylindrical, and is in turn a quasi-convex hierarchy for $G$.
Let $A$ be some edge group of the first splitting in this hierarchy, so by construction the normal closure of $A$ is the kernel of the homomorphism $G\rightarrow\mathbb{Z}$ that sends $a$ to $1$ and $b$ to 0. One may verify that this is not finitely generated using Brown’s algorithm [undefh, Theorem 4.3]. However, the group $G$ does fibre — by the same algorithm the homomorphism $\chi:G\rightarrow\mathbb{Z}$ that sends $a$ to $1$ and $b$ to 1-1 has finitely generated kernel, which we will call $F$.
Consider now the intersection of $F$ with any $G$-conjugate $A^{g}$ of $A$. We must have that $A^{g}\not\subset F$ because the normal closure of $A$ is the infinitely generated kernel of a map to $\mathbb{Z}$, and so the quotient of $A^{g}$ by $F\cap A^{g}$ must be isomorphic to $\mathbb{Z}$. It follows that, since $A^{g}$ is free of rank $3$, $F\cap A^{g}$ cannot be finitely generated. The splitting of $F$ induced by the above hierarchy therefore does not have finitely generated edge groups, but will be acylindrical (and indeed finite height by Theorem 3.8).
While the group $F$ in Example 3.9 is a non-elementary free group, so naturally admits a free splitting, the above argument illustrates that acylindrical graphs (or hierarchies) of groups, even in the relatively tame case of one-relator groups, may not induce splittings with finitely generated edge groups on their subgroups. However, we may still deal with such subgroups using methods from the proof of Theorem 3.3. We have the following corollary, which is a reformulation of Corollary 1.6.
@@
Let $G$ be a hyperbolic group, and let $H$ be a hyperbolic subgroup of $G$.
  1. (a)
    If $G$ admits an action on a tree $T$ that is a quasi-convex splitting over a locally quasi-convex group, and $H$ the induced action of $H$ on $T$ is non-trivial, then $H$ admits a non-trivial quasi-convex splitting that dominates the minimal subtree of the induced action of $H$ on $T$.
  2. (b)
    If $G$ admits a quasi-convex hierarchy with locally quasi-convex edge stabilisers then $H$ admits a quasi-convex hierarchy.
@@
We begin with part a. Let $T$ be a tree on which $G$ acts such that the action of $G$ on $T$ is a quasi-convex splitting, and let $T_{H}$ be the minimal subtree of the induced action of $H$ on $T$ guaranteed by Lemma 2.1. By Theorem 2.23 there exists some tree $T^{\prime}$ on which $H$ acts with finitely generated edge stabilisers such that $T^{\prime}$ dominates $T_{H}$ along the $H$-equivariant epimorphism $\phi:T^{\prime}\rightarrow T_{H}$. By Lemma 2.1 the action of $H$ on $T^{\prime}$ has a unique minimal subtree $T^{\prime}_{H}$ with finitely many orbits of edges, and the image of the restriction of $\phi$ to $T^{\prime}_{H}$ is a fixed subtree of $T_{H}$. This restriction must therefore be surjective by the fact that the action of $H$ on $T_{H}$ was minimal, and thus the action of $H$ on $T^{\prime}_{H}$ dominates the action of $G$ on $T$. As in the proof of Theorem 3.3, fix some $e\in E(T_{H})$ and consider the set of edges $S_{e}=\phi^{-1}(e)\subseteq E(T_{H}^{\prime})$, which we may equip with a metric given by considering the distance between their midpoints in $T_{H}^{\prime}$. By equivariance, $H_{e}=\operatorname{PStab}_{H}(e)$ acts on $S_{e}$ by isometry and this action has the properties analogous to property (1) in the proof of Theorem 3.3(a). In particular, the $H$-stabiliser $\operatorname{PStab}_{H}(f)$ of an element $f\in S_{e}$ with respect to the $H$-action on $T^{\prime}_{H}$ agrees with the $H_{e}$-stabiliser $\operatorname{PStab}_{H_{e}}(f)$ with respect to the $H_{e}$-action on $S_{e}$.
The $H$-stabiliser of each element of $f\in E(T_{H}^{\prime})$ is therefore a finitely generated subgroup of $H\cap G_{\phi(e)}$, where $G_{\phi(e)}=\operatorname{PStab}_{G}(e)$ is the $G$-stabiliser of $e$ with respect to the $G$-action on $T$. Furthermore, by assumption on the action of $G$ on $T$, $G_{e}$ is a locally quasi-convex quasi-convex subgroup of $G$, so every finitely generated subgroup of $G_{e}$ is a quasi-convex subgroup of $G$ by transitivity of quasi-convexity. It therefore follows that each stabiliser of the action of $H$ on $T_{H}^{\prime}$ is a quasi-convex subgroup of $G$, so is a quasi-convex subgroup of each hyperbolic subgroup of $G$. In particular, each stabiliser of the action of $H$ on $T_{H}^{\prime}$ is a quasi-convex subgroup of $H$ as required.
We now proceed with part b. Let $n$ a non-negative integer, and for all $0\leq i<n$ let $J_{i}$ be a finite indexing set with $J_{n}=\{1\}$. Let $\{G_{i,j}\}_{0\leq i\leq n,j\in J_{i}}$ be the set of groups in a quasi-convex hierarchy decomposition for $G$ of length $n$, as in Definition 2.3. We may assume without loss of generality that $H$ is not isomorphic to a subgroup of any element of $\{G_{i,j}\}_{0\leq i\leq n,j\in J_{i}}$ other than $G_{n,0}$, and proceed by induction on $n$. First assume $n=0$. Then $G$ is finite, and so $H$ is finite and admits a quasi-convex hierarchy of length 0 as required.
@@
Now assume that the above result holds for all hyperbolic subgroups of hyperbolic groups $G$ admitting a quasi-convex hierarchy of length at most k1k-1 with locally quasi-convex edge stabilisers, and assume that $n=k$. By property (H3) in Definition 2.3, $G\cong G_{n,1}$ admits an action on a tree $T$ with quasi-convex edge stabilisers and with vertex stabilisers of the form $G_{n-1,k}$ for some $k\in J_{n-1}$, and by assumption $H$ is not isomorphic to a subgroup of any vertex stabiliser of this action, so $H$ acts non-trivially on $T$. We may therefore apply part a of this corollary to see that $H$ admits a quasi-convex splitting that dominates the minimal subtree of the induced action of $H$ on $T$. The vertex stabilisers of this splitting will be subgroups of the vertex stabilisers of the action of $G$ on $T$ as above, and will be quasi-convex in $H$ by Lemma 2.17 and thus hyperbolic by Lemma 2.19(a). All vertex stabilisers of the action of $G$ on $T$ are of the form $G_{n-1,k}$ for some $k\in J_{n-1}$, and thus admit quasi-convex hierarchies of length at most n1n-1, and so it follows by the inductive hypothesis that each stabiliser of the action of $H$ on $T_{H}^{\prime}$ admits a quasi-convex hierarchy. It follows that $H$ must admit a quasi-convex hierarchy as required. ∎
The following corollary is a reformulation of Corollary 1.7, and therefore answers Question 1.3 in the affirmative for all finitely generated subgroups of one-relator groups with acylindrical hierarchies.
Corollary 3.11.
Let $H$ be a finitely generated subgroup of a one-relator group $G$ such that $G$ admits an acylindrical Magnus hierarchy $G_{0}\hookrightarrow G_{1}\hookrightarrow\cdots\hookrightarrow G_{n}\cong G$. Then $H$ admits a quasi-convex hierarchy. In particular the group $H$ is acylindrically arboreal if and only if it admits a non-elementary quasi-convex splitting.
Proof.
First, we observe that by Theorem 3.8 $G$ is hyperbolic and $G_{0}\hookrightarrow G_{1}\hookrightarrow\cdots\hookrightarrow G_{n}\cong G$ is a quasi-convex hierarchy whose edge stabilisers are free, and therefore locally quasi-convex by Halls’ theorem [undefn]. Furthermore, by [undefo, Corollary 9] $G$ has virtual cohomological dimension at most two, so there exists some finite index subgroup $G^{\prime}\leq G$ such that $\operatorname{cd}_{\mathbb{Z}}(G^{\prime})\leq 2$. By coherence of one-relator groups[undefq, Theorem 1.1] $H$ will be finitely presented, and $H^{\prime}=H\cap G^{\prime}$ is a finite index subgroup of $H$ so must also be finitely presented. It follows by Gersten’s theorem [undefl, Theorem 5.4] that $H^{\prime}$ is hyperbolic, and so $H$ must also be hyperbolic. We are therefore able to apply Corollary 1.6 to see that $H$ admits a quasi-convex hierarchy as required.
@@
Now assume that $H$ is acylindrically arboreal. It is therefore not virtually cyclic by Theorem 2.6, and so the hierarchy given by Corollary 1.6 in the first part of this proof will not be trivial. Taking the first stage of this hierarchy therefore gives the required non-trivial finite-height splitting with finitely generated edge stabilisers.
@@ 3.4. Relaxing The Conditions on Stabilisers
The group $G_{1}$ is recursively presented, so there exists a finitely presented group $G_{2}$ such that $G_{1}\leq G_{2}$ is a malnormal subgroup by a result of Wagner [undefah, Theorem 1.1]. Define $G$ to be the trivial HNN-extension of $G_{2}$ along $\langle t_{0}\rangle$, so $G=\langle G_{2},t\mid[t,t_{0}]\rangle$.
@@
[…]
so $)(H_{1}*_{\langle t_{0}\rangle}\mathbb{Z}^{2})\leq\operatorname{PStab}_{G}(\phi(\tilde{v}$)). By definition of $T$ this will hold only when $\phi(\tilde{v})=v$. It follows that $(H_{1}*_{\langle t_{0}\rangle}\mathbb{Z}^{2})\leq\operatorname{PStab}_{G}(\tilde{v})\leq(H_{1}*_{\langle t_{0}\rangle}\mathbb{Z}^{2})$, so $\operatorname{PStab}_{G}(\tilde{v})=(H_{1}*_{\langle t_{0}\rangle}\mathbb{Z}^{2})$ as required, and that at least one vertex stabiliser of the action of $G$ on $T^{\prime}$ is infinitely generated. The result then follows by Lemma 2.2. ∎
@@
The Dunwoody–Sageev resolution applied to the acylindrical splitting of the group $G$ above will not preserve acylindricity.
@@
which is acylindrical due to the malnormality of $G_{1}$ in $G_{2}$. We do however believe that there will be an example of a finitely presented group that acts acylindrically on a tree but not with finitely generated edge stabilisers, and finish this paper by restating Question 1.10.
Question 3.14.